OpenStudy (anonymous):

expand and simplify (2^x+5)^2

OpenStudy (turingtest):

\[(2^x+5)^2=(2^x+5)(2^x+5)=2^{2x}+10(2^x)+25\]\[=2^{2x}+5(2^{x+1})+25\]

OpenStudy (anonymous):

why is it 5(2^x+1)?

OpenStudy (turingtest):

because\[10\times2^x=5\times2\times2^x=5\times(2\times2^x)=5(2^{x+1})\]in the same way that\[x(x^n)=x^{n+1}\]I hope that helps...